OpenStudy (anonymous):

How do I find the x-intercept of y=x^3+5x^2-4x-20?

OpenStudy (kohai):

You factor and then solve for x for each of them

OpenStudy (kohai):

In this case it factors to (x+2)(x-2)(x+5) So you do it like this, x+2=0 x-2=0 x+5=0 x=-2, 2, -5